# distributive quandle and detecting causality

### v2

@@ Abstract
I investigated the capability of a specific type of quandle, quandle whose operation satisfying that $(a_{1}*b_{1})*(a_{2}*b_{2})=(a_{1}*a_{2})*(b_{1}*b_{2})$, to detect causality in (2+1)-dimensional globally hyperbolic spacetime by determining if they can distinguished the connected sum of two Hopf links with an infinite series of relevant three-component links constructed by Allen and Swenberg in 2020, who suggested that any link invariant must be able to distinguish those links for them to detect causality in the given setting. I show that these quandles fail to do so as long as $a\sim b\Leftrightarrow a*b=a$ defines an equivalence relation. The Alexander quandles is an example that this result can apply to.
@@ 1 Introduction
Section 2 of this paper introduced link invariant generated by quandle, an algebraic structure, and specifically for a specific type of quandle whose operation satisfy the property $(a_{1}*b_{1})*(a_{2}*b_{2})=(a_{1}*a_{2})*(b_{1}*b_{2})$ for all $a_{1},b-1,a_{2},b_{2}$ in the set, which I call the distributive quandle. In section 3-4, this paper investigate the capability of distributive quandles for distinguishing $L_{2H}$ and Allen-Swenberg links. I show that as long as the relation define on a distributive quandle by $a~{}b\Leftrightarrow a*b=a$ is an equivalence relation, then it cannot distinguish $L_{2H}$ and any of Allen-Swenberg link in the series, thus fail to capture causality in the spacetime. Since Alexander quandles satisfy the restriction on qunadle in this result, an application of this result is that all Alexander quandle fail to detect causality in the given spacetime.
@@ 3. $(x*y)*z=(x*z)*(y*z)$
definition 2.6 distributive quandle
Consider the type of quandle $Q(*)$ with the property that $\forall x_{1},y_{2},x_{2},y_{2}\in Q$, they satisfy that $(x_{1}*y_{2})*(x_{2}*y_{2})=(x_{1}*x_{2})*(y_{1}*y_{2})$. Let’s call this type of quandle the distributive quandle. Let $*^{-1}$ denote its inverse operation. It can be proved that a distributive quandle satisfies the following 9 properties, for all $x,y,z,x_{1},x_{2},y_{1},y_{2}\in Q$:
@@
An Alexander quandle is a module $M$ over $\mathbb{Z}[t^{\pm}]$ under the the operation $a*b=ta+(1-t)b$. Since we need finite quandle for quandle coloring link invariant, we consider finite Alexnader quandle which is in the form $Q=\mathbb{Z}_{n}[t^{\pm}]/(h(t))$ where $h$ is a monic polynomial in $t$ (Nelson[5]). It is a distributive quandle because $\forall a_{1},b_{1},a_{2},b_{2}\in Q$,
@@ 2.1 Example: coloring on the trefoil knot
3 distributive quandle & detecting causality
Allen and Swenberg[2] investigated whether there are link invariants can detect causality in $X$, a (2+1)-dimensional globally hyperbolic spacetime with a Cauchy surface $\sigma$ whose universal cover is homomorphic to $R^{2}$. They constructed an infinite series of three-component links (Allen-Swenberg links), and suggested that such link invariant needs to be able to distinguish between these links and connected sum of two Hopf ($L_{2H})$. (see figures of the links below). They found that Conway polynomial fail to do so. I would show the coloring invariant and its enhanced polynomial from any distributive quandle with $a\sim b\Leftrightarrow a*b=a$ defining an equivalence relation also can’t differentiate $L_{2H}$ from the Allen-Swenberg links.
@@
Let $Q(*)$ be a distributive quandle (definition 2.6) such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*b=a$ is an equivalence equation.
@@
This means any distributive quandle with the equivalence relation defined above is unable to distinguish between the any of Allen-Swenberg links and the connected sum of Hopf link by their coloring invariant or the enhanced version, therefore fail to detect causality in the given setting.
Example: Alexnader quandle
All finite Alexander quandle defined in section 2.7 satisfy the property of $Q(*)$ defined above. Section 2.7 show it is distributive. The relation $\sim$ defined on a finite Alexander quandle $A=\mathbb{Z}_{n}[t^{\pm}]/(h(t))$ is
@@
$(t-1)a=(t-1)b~{}(mod~{}h(t))$, which is clearly an equivalence relation. So proposition 3.1 imply that Alexnader quandle are unable to detect causality in the given spacetime.
@@
First we prove some properties of distributive quandle which will be used later. In some calculation process below, $=_{(i)}$ or $=_{k}$ means that the right side of the equation are derived by applying property $(i)$ in section 2.6 or property k in this section that is already proved. Let $Q(*)$ be a distributive quandle. Then
@@
Proof: A easy way to proof this is noticing that the orignal and inverse operation * and $*^{-1}$ is completely symmetric in a distributive quandle. This is because one can check that replacing all * by $*^{-1}$ and all $*^{-1}$ by * in any one of properties in section 2.6 gives another property that is alsp satisfied by $Q(*)$. So if an equation for operation in $Q(*)$ holds, than the equation derived by switching between all * and $*^{-1}$ in the orignal equation is also valid. By doing it to 1 we get $[(a*^{-1}x)*y]*^{-1}z=[(a*^{-1}z)*y]*^{-1}x$, which is exactly 2, which complet the proof.
@@
Let $Q(*)$ be a distributive quandle such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*c=a$ is an equivalence equation. For the connected sum of two Hopf links, the system of equations associated to the coloring space comes from only 4 crossings, so it can be calculated relatively easily. Consider the following label by a valid coloring on $L_{2H}$ by $Q$:
@@ 4.2 Lemma 1
Claim: when applying coloring from any distributive quandle $Q(*)$, the system of equations with the operation from $Q(*)$ given by this subpart link diagram force $n=d$ and $m=e$.
@@
Let $Q(*)$ be an arbitrary distributive quandle. Let’s first analysis the coloring of $Q(*)$ on the following two structures, called structure ① and ②,respectively. In both structures by labeling 4 of the arcs exposed outside with $x,y,n,m$ representing some element in $Q(*)$, the other 4 out-exposed aces labeled by $x^{\prime},y^{\prime},n^{\prime},m^{\prime}$ can be expressed by a string of quandle operations between $x,y,n,m$. By using the properties of distributive quandle given in section 2.6, these expressions can be further simplified to the form of some functions $f_{pq},~{}p,q\in\{x,y,n,m\}$ defined in definition 4.1 acting on one of $x,y,n,m$ The result is shown in the figure below and the calculation process is shown below.
@@
Pluging in $j=b,~{}i=a$, the above equations gives $d=f_{ji}\circ f_{yx}\circ f_{ij}\circ f_{xy}(n)$ and $e=f_{ji}\circ f_{yx}\circ f_{ij}\circ f_{xy}(m)$.
@@
Let $Q(*)$ be a distributive quandle such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*c=a$ is an equivalence equation. Then referring to figure 4.3, its coloring space on the first Allen-Swenberg link $A_{1}$ is the solution set $\{x1=x2=x3=x4=c_{1},y1=y2=c_{2},~{}others=c_{3}\}$, where $c_{1},c_{2},c_{3}\in Q$ with the condition that $c_{1}\sim c_{2}\sim c_{3}$$others"$ refer to all line segment in figure 3 other than $x1,x2,x2,x4,y1,y2$.
@@
In the oriented $A_{k}$ digram, we can apply lemma 1 to label all arcs in the middle component outside each $T_{i}$ by $m_{j}$ where $0\geq j\leq 2n$:
@@
By section 4.1,4.4 and 4.5, if $Q(*)$ is a distributive quandle such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*c=a$ is an equivalence equation, the solutions sets correspond to the coloring space for both connected sum of two Hopf links $L_{2H}$ and general Allen-Swenberg links $A_{k}$ are in the form $\{x=c_{1},y=c_{2},z=c_{3}~{}\forall x\in X,y\in Y,z\in Z~{}|~{}c_{1}\sim c_{2}\sim c_{3}\in Q(*)\}$, where $X,Y,Z$ is a partition of the correspond fundamental quandle of the link $Q(L)$. This means the coloring space both $L_{2H}$ and any $A_{k}$ should be $\{f:Q(L)\to Q(*)~{}|~{}f(X)=\{c_{1}\},f(Y)=\{c_{2}\},f(Z)=\{c_{3}\},c_{1}\sim c_{2}\sim c_{3}\}$.

### v3

# medial quandle and detecting causality

@@ Abstract
I investigated the capability of medial quandle, quandle whose operation satisfying that $(a_{1}*b_{1})*(a_{2}*b_{2})=(a_{1}*a_{2})*(b_{1}*b_{2})$, to detect causality in (2+1)-dimensional globally hyperbolic spacetime by determining if they can distinguished the connected sum of two Hopf links with an infinite series of relevant three-component links constructed by Allen and Swenberg in 2020, who suggested that any link invariant must be able to distinguish those links for them to detect causality in the given setting. I show that these quandles fail to do so as long as $a\sim b\Leftrightarrow a*b=a$ defines an equivalence relation. The Alexander quandles is an example that this result can apply to.
@@ 1 Introduction
Section 2 of this paper introduced link invariant generated by quandle, an algebraic structure, and specifically for medial quandle. In section 3-4, this paper investigate the capability of medial quandle for distinguishing $L_{2H}$ and Allen-Swenberg links. I show that as long as the relation define on a medial quandle by $a~{}b\Leftrightarrow a*b=a$ is an equivalence relation, then it cannot distinguish $L_{2H}$ and any of Allen-Swenberg link in the series, thus fail to capture causality in the spacetime. Since Alexander quandles satisfy the restriction on qunadle in this result, one application of this result is that all Alexander quandle fail to detect causality in the given spacetime.
@@ 3. $(x*y)*z=(x*z)*(y*z)$
definition 2.6 medial quandle
Medial quandle is the type of quandle $Q(*)$ such that $\forall x_{1},y_{2},x_{2},y_{2}\in Q$, they satisfy that $(x_{1}*y_{2})*(x_{2}*y_{2})=(x_{1}*x_{2})*(y_{1}*y_{2})$. Let $*^{-1}$ denote its inverse operation. It can be proved that a medial quandle satisfies the following 9 properties, for all $x,y,z,x_{1},x_{2},y_{1},y_{2}\in Q$:
@@
An Alexander quandle is a module $M$ over $\mathbb{Z}[t^{\pm}]$ under the the operation $a*b=ta+(1-t)b$. Since we need finite quandle for quandle coloring link invariant, we consider finite Alexnader quandle which is in the form $Q=\mathbb{Z}_{n}[t^{\pm}]/(h(t))$ where $h$ is a monic polynomial in $t$ (Nelson[5]). It is a medial quandle because $\forall a_{1},b_{1},a_{2},b_{2}\in Q$,
@@ 2.1 Example: coloring on the trefoil knot
3 medial quandle & detecting causality
Allen and Swenberg[2] investigated whether there are link invariants can detect causality in $X$, a (2+1)-dimensional globally hyperbolic spacetime with a Cauchy surface $\sigma$ whose universal cover is homomorphic to $R^{2}$. They constructed an infinite series of three-component links (Allen-Swenberg links), and suggested that such link invariant needs to be able to distinguish between these links and connected sum of two Hopf ($L_{2H})$. (see figures of the links below). They found that Conway polynomial fail to do so. I would show the coloring invariant and its enhanced polynomial from any meidal quandle with $a\sim b\Leftrightarrow a*b=a$ defining an equivalence relation also can’t differentiate $L_{2H}$ from the Allen-Swenberg links.
@@
Let $Q(*)$ be a medial quandle (definition 2.6) such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*b=a$ is an equivalence equation.
@@
This means any medial quandle with the equivalence relation defined above is unable to distinguish between the any of Allen-Swenberg links and the connected sum of Hopf link by their coloring invariant or the enhanced version, therefore fail to detect causality in the given setting.
Example: Alexander quandle
All finite Alexander quandle defined in section 2.7 satisfy the property of $Q(*)$ defined above. Section 2.7 show it is medial. The relation $\sim$ defined on a finite Alexander quandle $A=\mathbb{Z}_{n}[t^{\pm}]/(h(t))$ is
@@
$(t-1)a=(t-1)b~{}(mod~{}h(t))$, which is clearly an equivalence relation. So proposition 3.1 imply that Alexander quandle are unable to detect causality in the given spacetime.
@@
First we prove some properties of medial quandle which will be used later. In some calculation process below, $=_{(i)}$ or $=_{k}$ means that the right side of the equation are derived by applying property $(i)$ in section 2.6 or property k in this section that is already proved. Let $Q(*)$ be a medial quandle. Then
@@
Proof: A easy way to proof this is noticing that the original and inverse operation * and $*^{-1}$ is completely symmetric in a medial quandle. This is because one can check that replacing all * by $*^{-1}$ and all $*^{-1}$ by * in any one of properties in section 2.6 gives another property that is also satisfied by $Q(*)$. So if an equation for operation in $Q(*)$ holds, than the equation derived by switching between all * and $*^{-1}$ in the original equation is also valid. By doing it to 1 we get $[(a*^{-1}x)*y]*^{-1}z=[(a*^{-1}z)*y]*^{-1}x$, which is exactly 2, which completes the proof.
@@
Let $Q(*)$ be a medial quandle such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*c=a$ is an equivalence equation. For the connected sum of two Hopf links, the system of equations associated to the coloring space comes from only 4 crossings, so it can be calculated relatively easily. Consider the following label by a valid coloring on $L_{2H}$ by $Q$:
@@ 4.2 Lemma 1
Claim: when applying coloring from any medial quandle $Q(*)$, the system of equations with the operation from $Q(*)$ given by this subpart link diagram force $n=d$ and $m=e$.
@@
Let $Q(*)$ be an arbitrary medial quandle. Let’s first analysis the coloring of $Q(*)$ on the following two structures, called structure ① and ②,respectively. In both structures by labeling 4 of the arcs exposed outside with $x,y,n,m$ representing some element in $Q(*)$, the other 4 out-exposed aces labeled by $x^{\prime},y^{\prime},n^{\prime},m^{\prime}$ can be expressed by a string of quandle operations between $x,y,n,m$. By using the properties of medial quandle given in section 2.6, these expressions can be further simplified to the form of some functions $f_{pq},~{}p,q\in\{x,y,n,m\}$ defined in definition 4.1 acting on one of $x,y,n,m$ The result is shown in figure 4.2 below and the calculation process is shown below.
@@
Plugging in $j=b,~{}i=a$, the above equations gives $d=f_{ji}\circ f_{yx}\circ f_{ij}\circ f_{xy}(n)$ and $e=f_{ji}\circ f_{yx}\circ f_{ij}\circ f_{xy}(m)$.
@@
Let $Q(*)$ be a medial quandle such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*c=a$ is an equivalence equation. Then referring to figure 4.3, its coloring space on the first Allen-Swenberg link $A_{1}$ is the solution set $\{x1=x2=x3=x4=c_{1},y1=y2=c_{2},~{}others=c_{3}\}$, where $c_{1},c_{2},c_{3}\in Q$ with the condition that $c_{1}\sim c_{2}\sim c_{3}$$others"$ refer to all line segment in figure 3 other than $x1,x2,x2,x4,y1,y2$.
@@
In the oriented $A_{k}$ diagram, we can apply lemma 1 to label all arcs in the middle component outside each $T_{i}$ by $m_{j}$ where $0\geq j\leq 2n$:
@@
By section 4.1,4.4 and 4.5, if $Q(*)$ is a medial quandle such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*c=a$ is an equivalence equation, the solutions sets correspond to the coloring space for both connected sum of two Hopf links $L_{2H}$ and general Allen-Swenberg links $A_{k}$ are in the form $\{x=c_{1},y=c_{2},z=c_{3}~{}\forall x\in X,y\in Y,z\in Z~{}|~{}c_{1}\sim c_{2}\sim c_{3}\in Q(*)\}$, where $X,Y,Z$ is a partition of the correspond arcs set of the link $Q(L)$. This means the coloring space both $L_{2H}$ and any $A_{k}$ should be $\{f:Q(L)\to Q(*)~{}|~{}f(X)=\{c_{1}\},f(Y)=\{c_{2}\},f(Z)=\{c_{3}\},c_{1}\sim c_{2}\sim c_{3}\}$.